The position of a particle is given by x(t)=4.0t^2+0.5t^3m



a. Find the instantaneous velocity at t=5.0 to 3.0sec





b. What will be the average velocity fist =5.0 and 6.0

1
Expert's answer
2022-01-12T08:28:29-0500

(a) The instantaneous velocity is the derivative of position by time:


"v(t)=8t+1.5t^2,\\
v(5)=77.5\text{ m/s},\\
v(3)=37.5\text{ m/s}."


(b) The average velocity is


"v_\text{av}=\dfrac{x(6)-x(5)}{6-5}=75.92\text{ m/s}."